Use the properties of logarithms to solve the equation for x.
ln(8x − 13) − ln(x − 3) = ln 11
Thanks!!

Answer by stanbon(75887)

Use the properties of logarithms to solve the equation for x.
ln(8x − 13) − ln(x − 3) = ln 11
---------
ln[(8x-13)/(x-3)] = ln(11)
--------
(8x-13)/(x-3) = 11
8x-13 = 11x-33
3x = 20
x = 20/3
------------------
Cheers,
Stan H.
===================